OpenStudy (anonymous):

Given the formula, solve for the indicated variable. F= 9/5C +32 for C

OpenStudy (diyadiya):

\[F= \frac{9}{5}C+32\] subtract 32 from both sides \[F-32= \frac{9}{5}C+32-32\]\[F-32= \frac{9}{5}C\]Multiply both sides by 5\[5(F-32)=5 \times \frac{9}{5}C =5(F-32)=\cancel{5} \times \frac{9}{\cancel{5}}C\]\[5(F-32)=9C\]Divide both sides by 9\[\frac{5(F-32)}{9}= \frac{9C}{9}\]\[\frac{5(F-32)}{9}= \frac{\cancel{9}C}{\cancel{9}}\]\[C=\frac{5(F-32)}{9} or C= \frac{5}{9} (F-32)\]

OpenStudy (anonymous):

thanks